OpenStudy (anonymous):

does 2157 divide by 3?

OpenStudy (ivettef365):

let see 21/3 = 7 and 57/3 = 19 so it definitely divides by 3

OpenStudy (mathstudent55):

Add the digits of 2157 together. If the sum divides by 3, then 2157 divides by 3.

OpenStudy (anonymous):

Any number is divisible by 3 iff the sum of all the numbers is divisible Here 2+1+5+7=15 As 15 is divisible by 3 So the given number is also divisible by 3